# SDP bounds on quantum codes

### v1

@@ Appendix A Dual programs
By eliminiating the slack variable $Y$ and redefining $y_{i}$ with $-y_{i}$, one obtains
@@ A.1. Dual of the Lovász bound
We use this to write the dual program of the feasibility version of the symmetry reduced quantum Lovász number for self-dual codes,
@@
$y^{t,p}_{i,j}=0\hskip 80.00012pt\text{if}\quad i,j\neq 0\,,\quad t-p\quad\text{is even }\quad$
@@
where $Y^{(a,k)}$ with $0\leq a\leq k\leq n+a-k$ are real square matrices of size $(n+a-2k)$.
@@
where $\alpha(i,j,t,p,a,k)$ is given by Eq. (6.3). Note that the condition $x^{t,p}_{i,j}=0$ if $0<i+j-t-p<\delta$ is equivalent to the pure code condition from Eq. (141). This can be seen from that fact that when $j=0$, it simplifies to $x^{0,0}_{i,0}=x^{i,i}_{i,i}=0$ if $0<i<\delta$. But $x_{i,i}^{i,i}=0$ implies that $x_{i,j}^{t,p}=0$ for any $j,t,p$: Consider $x\in{\boldsymbol{4}}^{n}$ with $s(x)=i$. Then $\tilde{\Gamma}_{xx}=x^{i,i}_{i,i}=0$. The constraint $\tilde{\Gamma}\succeq 0$ implies that $\tilde{\Gamma}_{xy}=x^{t,p}_{i,j}=0$ for any $y$ with $|s(y)|=j$, $\lvert s(x)\cap s(y)\rvert=t$ and |$\lvert\{m\,\mid\,x_{m}=y_{m}\neq 0\}\rvert=p$. This is due to the fact that the $2\times 2$ minor of $\tilde{\Gamma}$ involving $x$ and $y$ must be positive-semidefinite, and thus if $\tilde{\Gamma}_{xx}=0$ then also $\tilde{\Gamma}_{xy}=0$.
@@
=$\mathcal{R}=\big{\{}(i,j,t,p,a,k)\,\big{|}\,\,$ $(i,j,t,p)\in\mathcal{I}(4,n)\,,\quad 0\leq k\leq\min(i,j)\,,$
@@
Otherwise, it tends to $\infty$- ∞. Note that conditions (2) and (5) appearing in Eq. (167),
@@
00 0 $=y_{i,i}^{i,i}+\gamma_{i,0}^{0,0}w+2y_{i,0}^{0,0}\quad$ if $\delta\leq i\leq n$
(169) 00 0 $=y^{t,p}_{i,j}$ if $i,j\neq 0,\quad(t-p)\,\,\text{even}\,,\quad i+j-t-p\geq\delta\,.$
@@
$y^{t,p}_{i,j}=0\hskip 85.00012pt\text{if}\quad i,j\neq 0\,,\quad t-p\quad\text{is even }\quad$
@@ A.2. Dual of the symmetry-reduced SDP
Here the dual variables are real matrices $Y^{(a,k)}$ of size $n+a-2k$ with $0\leq a\leq k\leq n+a-k$ and $Q_{i},C_{i}\in\mathds{R}$ with $0\leq i\leq n$.
@@
permutation $\hskip 73.97733pt\text{permutation of}\,\,(i^{\prime},j^{\prime},i^{\prime}+j^{\prime}-t^{\prime}-p^{\prime})\quad\text{with}\quad i,j,i+j-t-p\neq 0\,,$
@@
Similarly to the previous section, we now construct the Lagrangian. Associate with each equality constraint a dual variable ${u},{v_{i}},r^{t,p}_{i,j},s^{t,t^{\prime},p,p^{\prime}}_{i,i^{\prime},j,j^{\prime}},Q_{i},C_{i}\in\mathds{R}$ and with each semidefinite constraint a dual semidefinite variable $Y^{(a,k)}\succeq 0$ and the objective function equals to 00. Then
@@
From the definition of $f^{t,p}_{i,j}$ follows a first constraint: for $(a,b,c,d)\in\mathcal{I}(4,n)$ with $(c-d)$ even and $a,b,a+b-c-d~{}\neq~{}0$,
@@
Otherwise, it tends to $\infty$- ∞. Similarly to Eq. (A.1), condition (5) above,
@@
(182) 00$\quad$ 0 $=\gamma^{0,0}_{i,0}(KD_{i}-C_{i})+\left(\frac{2^{n}}{K}-2\right)\gamma^{0,0}_{i,0}Q_{i}-2y^{0,0}_{i,0}-\gamma^{i,i}_{i,i}Q_{0}-y^{i,i}_{i,i}\quad\text{if}\quad 0<i<\delta\,,$
(183) 00$\quad$ 0 $=\gamma^{0,0}_{i,0}KD_{i}+\left(\frac{2^{n}}{K}-2\right)\gamma^{0,0}_{i,0}Q_{i}-2y^{0,0}_{i,0}-\gamma^{i,i}_{i,i}Q_{0}-y^{i,i}_{i,i}\quad\quad\quad\quad\text{if}\quad\delta\leq i\leq n\,.$
@@
Finally, one can check from Eq. (77) that $\gamma^{t,p}_{ij}=\gamma^{t^{\prime},p^{\prime}}_{i^{\prime}j^{\prime}}$ if $t-p=t^{\prime}-p^{\prime}$ is even and $(i,j,i+j-t-p)$ is a permutation of $(i^{\prime},j^{\prime},i^{\prime}+j^{\prime}-t^{\prime}-p^{\prime})$. In particular, $\gamma^{0,0}_{i,0}=\gamma^{i,i}_{i,i}$. Thus, by mapping the variable $y^{t,p}_{ij}\mapsto\gamma^{t,p}_{i,j}y^{t,p}_{ij}$, the constraints in the SDP simplify to Eq. (26) since $\gamma^{t,p}_{ij}$ factorizes. This ends the proof. ∎
@@ Appendix B Infeasibility certificate $(\!(7,1,4)\!)_{2}$ code
The following matrices $Y^{(a,k)}$ are a solution of the dual program (9.1) for $n=7$ and $\delta=4$ with $\alpha=0.58>0$ and a maximum constraint violation of $\mathcal{O}(10^{-16})$. Note that the first constraint from Eq. (25) allows to determine the value from $w$ using $Y^{(a,k)}$ since $w=-y^{i,i}_{i,i}-y^{0,0}_{i,0}$ if $\delta\leq i\leq n$. In particular, one can choose $w=-y^{n,n}_{n,n}-y^{0,0}_{n,0}$. The martices thus provide an infeasibility certificate and show that a $(\!(7,1,4)\!)_{2}$ quantum code does not exist.

### v2

@@ Appendix A Dual programs
By eliminiating the slack variable $Y$ and redefining $y_{i}$ with $y_{i}$, one obtains
@@ A.1. Dual of the Lovász bound
We use this to write the dual program of the feasibility version of the symmetry reduced quantum Lovász theta number for self-dual codes,
@@
[…]
@@
where $Y^{(a,k)}$ with $0\leq a\leq k\leq n+a-k$ are real square matrices of size (n+a2k)(n+a-2k).
@@
where $\alpha(i,j,t,p,a,k)$ is given by Eq. (6.3). Note that the condition $x^{t,p}_{i,j}=0$ if $0<i+j-t-p<\delta$ is equivalent to the pure code condition from Eq. (140). This can be seen from that fact that when $j=0$, it simplifies to $x^{0,0}_{i,0}=x^{i,i}_{i,i}=0$ if $0<i<\delta$. But $x_{i,i}^{i,i}=0$ implies that $x_{i,j}^{t,p}=0$ for any $j,t,p$: Consider $x\in{\boldsymbol{4}}^{n}$ with $s(x)=i$. Then $\tilde{\Gamma}_{xx}=x^{i,i}_{i,i}=0$. The constraint $\tilde{\Gamma}\succeq 0$ implies that $\tilde{\Gamma}_{xy}=x^{t,p}_{i,j}=0$ for any $y$ with $|s(y)|=j$, $\lvert s(x)\cap s(y)\rvert=t$ and $\lvert\{m\,\mid\,x_{m}=y_{m}\neq 0\}\rvert=p$. This is due to the fact that the $2\times 2$ minor of $\tilde{\Gamma}$ involving $x$ and $y$ must be positive-semidefinite, and thus if $\tilde{\Gamma}_{xx}=0$ then also $\tilde{\Gamma}_{xy}=0$.
@@
$\mathcal{R}=\big{\{}(i,j,t,p,a,k)\,\big{|}\,\,$ $(i,j,t,p)\in\mathcal{I}(4,n)\,,\quad 0\leq k\leq\min(i,j)\,$,
@@
Otherwise, it tends to -$\infty$. Note that conditions (2) and (5) appearing in Eq. (166),
@@
$0$ $=y_{i,i}^{i,i}+\gamma_{i,0}^{0,0}w+2y_{i,0}^{0,0}\quad$ if $\delta\leq i\leq n$
(168) $0$ $=y^{t,p}_{i,j}$ if $i,j\neq 0,\quad(t-p)\,\,\text{even}\,,\quad i+j-t-p\geq\delta\,.$
@@
[…]
@@ A.2. Dual of the symmetry-reduced SDP
Here the dual variables are real matrices $Y^{(a,k)}$ of size n+a2kn+a-2k with $0\leq a\leq k\leq n+a-k$ and $Q_{i},C_{i}\in\mathds{R}$ with $0\leq i\leq n$.
@@
$\hskip 73.97733pt\text{permutation of}\,\,(i^{\prime},j^{\prime},i^{\prime}+j^{\prime}-t^{\prime}-p^{\prime})\quad\text{with}\quad i,j,i+j-t-p\neq 0\,$,
@@
Similarly to the previous section, we now construct the Lagrangian. Associate with each equality constraint a dual variable ${u},{v_{i}},r^{t,p}_{i,j},s^{t,t^{\prime},p,p^{\prime}}_{i,i^{\prime},j,j^{\prime}},Q_{i},C_{i}\in\mathds{R}$ and with each semidefinite constraint a dual semidefinite variable $Y^{(a,k)}\succeq 0$ and the objective function equals to 0. Then
@@
From the definition of $f^{t,p}_{i,j}$ follows a first constraint: for $(a,b,c,d)\in\mathcal{I}(4,n)$ with (cd)(c-d) even and $a,b,a+b-c-d~\neq~0$,
@@
Otherwise, it tends to -$\infty$. Similarly to Eq. (A.1), condition (5) above,
@@
(181) $\quad 0$ $=\gamma^{0,0}_{i,0}(KD_{i}-C_{i})+\left(\frac{2^{n}}{K}-2\right)\gamma^{0,0}_{i,0}Q_{i}-2y^{0,0}_{i,0}-\gamma^{i,i}_{i,i}Q_{0}-y^{i,i}_{i,i}\quad\text{if}\quad 0<i<\delta\,$,
(182) $\quad 0$ $=\gamma^{0,0}_{i,0}KD_{i}+\left(\frac{2^{n}}{K}-2\right)\gamma^{0,0}_{i,0}Q_{i}-2y^{0,0}_{i,0}-\gamma^{i,i}_{i,i}Q_{0}-y^{i,i}_{i,i}\quad\quad\quad\quad\text{if}\quad\delta\leq i\leq n\,.$
@@
Finally, one can check from Eq. (79) that $\gamma^{t,p}_{ij}=\gamma^{t^{\prime},p^{\prime}}_{i^{\prime}j^{\prime}}$ if $t-p=t^{\prime}-p^{\prime}$ is even and (i,j,i+jtp)(i,j,i+j-t-p) is a permutation of $(i^{\prime},j^{\prime},i^{\prime}+j^{\prime}-t^{\prime}-p^{\prime})$. In particular, $\gamma^{0,0}_{i,0}=\gamma^{i,i}_{i,i}$. Thus, by mapping the variable $y^{t,p}_{ij}\mapsto\gamma^{t,p}_{i,j}y^{t,p}_{ij}$, the constraints in the SDP simplify to Eq. (27) since $\gamma^{t,p}_{ij}$ factorizes. This ends the proof. ∎
@@ Appendix B Infeasibility certificate $(\!(7,1,4)\!)_{2}$ code
The following matrices $Y^{(a,k)}$ are a solution of the dual program (9.1) for $n=7$ and $\delta=4$ with $\alpha=0.58>0$ and a maximum constraint violation of $\mathcal{O}(10^{-16})$. Note that the first constraint from Eq. (26) allows to determine the value from $w$ using $Y^{(a,k)}$ since $w=-y^{i,i}_{i,i}-y^{0,0}_{i,0}$ if $\delta\leq i\leq n$. In particular, one can choose $w=-y^{n,n}_{n,n}-y^{0,0}_{n,0}$. The matrices thus provide an infeasibility certificate and show that a $(\!(7,1,4)\!)_{2}$ quantum code does not exist.